# CLQCD Collaboration

### v1

@@ Abstract
The leading-order low-energy constant $\Delta_{\rm mix}$ in mixed-action chiral perturbation theory is calculated using $2+1+1$-flavor gauge ensembles with HISQ fermions and a tadpole-improved Symanzik gauge action at four lattice spacings $a\in[0.048,0.111]$ fm. By comparing our results to those from different actions and a $2+1$-flavor case, we find that the fermion action has the major impact, but the gauge action and charm quark loops also contribute significantly.
@@ II Setup and basic tests of the generated gauge ensembles
where the smeared link field $W_{\mu}(x)$ is constructed via a two-stage Fat7 smearing process with intermediate SU(3) projection, and $X_{\mu}(x)$ denotes links smeared with a single application of the ASQTAD smearing operator. The parameter provides improved dispersion relation for heavy quarks with bare mass $m$ by canceling tree-level $\mathcal{O}((am)^{4})$ errors.
@@
$S_{g}=\hat{\beta}\bigg$[ $\sum_{P}\left(1-\frac{1}{3}\mathrm{Re}\mathrm{Tr}(P)\right)$ (4)
@@
$a_{w_{0}}(\hat{\beta},\tilde{m}_{\pi},\tilde{m}_{\eta_{s}},\delta u_{0},\delta v_{0})=a(\hat{\beta})\big[1+$
@@
A comparison reveals that at a similar lattice spacing of $a\sim 0.11$ fm, the value of $\hat{\beta}$ is 1.1% larger for the 2+1+1 flavor ensembles than for the 2+1 flavor case. This indicates that the tadpole improved bare strong coupling $\alpha^{\rm tad}_{s}$ decreases by approximately 2.0% when charm sea quark effects are included. Using the same formulation, we find that $\alpha_{s}^{\rm tad}$ obtained with the 2+1 flavor HISQ action is 1.2% and 2.8% smaller when paired with the tree-level Symanzik or Iwasaki gauge action, respectively, compared to the result with the tadpole-improved Symanzik gauge action. In contrast, changing either the fermion action from SC to HI or the gauge action from Stad to S(1) leads to a much more significant shift in $\hat{\beta}$, corresponding to an $\sim 18$% change in $\alpha^{\rm tad}_{s}$.
@@ III Result of $\Delta_{\rm mix}$
Even-grid source (including $\tilde{L}^{3}/8$ points at all the spatial positions where the x/y/z indices are even) are used generate propagator for all the valence fermion actions to improve the statistics.
@@
Figure 4: The mixed action effects $\Delta_{\rm mix}$ at $m_{\pi}\sim 0.30$ GeV for various valence fermion actions (OV, SC and HC) and $\bar{\Delta}_{\rm mix,uni}$ for different valence fermion actions (OV+HC, OV+SC and HC+HC) on the 2+1 flavor DW+I, HI+I, HI+Stad, HI+S(0), and CLstout ensembles, and also 2+1+1 flavor HI+S(1), and HI+Stad ensembles. All the quantities are in unit of GeV2, and are interpolated to $a=0.11$ fm and $m_{\pi}\sim 300$ MeV to make a fair comparison.
@@
Comparing the 2+1 and 2+1+1 flavor HISQ ensembles with the same Stad gauge action, the inclusion of charm sea quarks appears to suppress $\Delta_{\text{mix}}$ by approximately 15%. However, the $\Delta_{\text{mix}}$ values for both of these cases are lower than that of the 2+1+1 flavor HISQ ensemble with the S(1) gauge action.
@@ IV Summary
Our central finding is the observation of a favorable $\mathcal{O}(a^{4})$ scaling for $\Delta_{\text{mix}}$ when valence overlap or clover fermions are used on gauge ensembles generated with 2+1+1 flavor HISQ sea fermions. This result is particularly significant as it was achieved using the same tadpole-improved Symanzik gauge action previously associated with a sizable $\Delta_{\text{mix}}$ in CLQCD clover fermion ensembles Zhao:2022ooq. This comparison directly verifies that the use of a dynamical fermion action with chiral symmetry is the dominant factor in suppressing mixed-action artifacts, irrespective of the gauge action. Furthermore, by comparing different action combinations at a fixed lattice spacing, we have quantified additional, secondary influences on $\Delta_{\text{mix}}$. We find that both the specific choice of improved gauge action and the inclusion of dynamical charm quark effects provide a further suppression of $\Delta_{\text{mix}}$.
@@ IV.1 Autocorrelation Analysis
Figure 5: Autocorrelation analysis across multiple lattice ensembles, depicting the bin-size dependence of variance for four fundamental quantities: the plaquette operator, topological charge, pion mass, and $\eta_{c}$ meson mass. The notation $\{n_{\mathrm{min}},n_{\mathrm{max}},\Delta n\}$ accompanying each observable specifies the minimal and maximal trajectory indices, along with the stride parameter employed for configuration subsampling to ensure statistical independence in our analysis.

### v2

@@ Abstract
The leading-order low-energy constant $\Delta_{\rm mix}$ in mixed-action chiral perturbation theory is calculated using $2+1+1$-flavor gauge ensembles with HISQ fermions and a tadpole-improved Symanzik gauge action at four lattice spacings $a\in[0.048,0.111]$ fm. By comparing our results to those from different actions and a $2+1$-flavor case, We find that the fermion action has the dominant impact, the gauge action has a secondary but measurable effect, and the contribution from charm quark loops is negligible within our current uncertainties.
@@ II Setup and basic tests of the generated gauge ensembles
where the smeared link field $W_{\mu}(x)$ is constructed via a two-stage Fat7 smearing process with intermediate SU(3) projection, and $X_{\mu}(x)$ denotes links smeared with a single application of the ASQTAD smearing operator. The parameter $\epsilon$ provides improved dispersion relation for heavy quarks with bare mass $m$ by canceling tree-level $\mathcal{O}((am)^{4})$ errors.
@@
$S_{g}=\hat{\beta}\bigg$[ $\sum_{P}\left(1-\frac{1}{3}\mathrm{Re}\mathrm{Tr}(P)\right)$
@@
$a_{w_{0}}(\hat{\beta},\tilde{m}_{\pi},\tilde{m}_{\eta_{s}},\delta u_{0})=a(\hat{\beta})\big[1+$
@@
A comparison reveals that at a similar lattice spacing of $a=0.1116$ fm, the value of $\hat{\beta}$ is 0.6% larger for the 2+1+1 flavor ensembles than for the 2+1 flavor case. This indicates that the tadpole improved bare strong coupling $\alpha^{\rm tad}_{s}$ decreases by approximately 1.1% when charm sea quark effects are included. Using the same formulation, we find that $\alpha_{s}^{\rm tad}$ obtained with the 2+1 flavor HISQ action is 1.2% and 2.8% smaller when paired with the tree-level Symanzik or Iwasaki gauge action, respectively, compared to the result with the tadpole-improved Symanzik gauge action. In contrast, changing either the fermion action from SC to HI or the gauge action from Stad to S(1) leads to a much more significant shift in $\hat{\beta}$, corresponding to an $\sim 18$% change in $\alpha^{\rm tad}_{s}$.
@@ III Result of $\Delta_{\rm mix}$
Even-grid source (including $\tilde{L}^{3}/8$ points at all the spatial positions where the x/y/z indices are even) are used generate propagator for all the valence fermion actions to improve the statistics. In practice, on each ensemble we compute quark propagators for the sea (HISQ) action and for the valence action at several valence quark masses $m_{v}$. We construct pseudoscalar two-point functions for the sea–sea (ss), valence–valence (vv), and valence–sea (vs) channels and extract $m_{\pi,ss}$, $m_{\pi,vv}(m_{v})$, and $m_{\pi,vs}(m_{v})$ from standard fits. We then interpolate in $m_{v}$ to the matching point where $m_{\pi,vv}=m_{\pi,ss}$, and evaluate $\Delta_{\mathrm{mix}}$ using Eq. (1) at this point. Since our sea action employs HISQ fermions, calculating $\Delta_{\text{mix}}$ necessitates an expansion of the Dirac indices for the corresponding HISQ quark propagator $S^{\text{HISQ}}$. For this purpose, we extend $S^{\text{HISQ}}$ into unitary matrices in the Dirac space, and multiply products of gamma matrices at each sink point $(x_{\text{snk}},y_{\text{snk}},z_{\text{snk}},t_{\text{snk}})$ of the even-grid source propagator, namely $S^{\text{Dirac}}=\gamma_{t}^{t_{\text{snk}}}\gamma_{z}^{z_{\text{snk}}}\gamma_{y}^{y_{\text{snk}}}\gamma_{x}^{x_{\text{snk}}}S^{\text{HISQ}}$.
@@
Figure 4: The mixed action effects $\Delta_{\rm mix}$ at $m_{\pi}\sim 0.30$ GeV for various valence fermion actions (OV, SC and HC) and $\bar{\Delta}_{\rm mix,uni}$ for different valence fermion actions (OV+HC, OV+SC and HC+HC) on the 2+1 flavor HI+I, HI+Stad, HI+S(0) ensembles, 2+1+1 flavor HI+Stad ensembles, and also includes the values from Ref. Zhao:2022ooq (shown as histograms in the shaded region) for comparison
. All the quantities are in unit of GeV2, and are interpolated to $a=0.11$ fm and $m_{\pi}\sim 300$ MeV to make a fair comparison.
@@
A comparison between the 2+1 and 2+1+1 flavor HISQ ensembles with the same Stad gauge action reveals that $\Delta_{\text{mix}}$ is consistent within statistical uncertainties, regardless of the inclusion of charm sea quarks, once the lattice spacing is rescaled to a common value assuming $a^{4}$ scaling. In contrast, both of these $\Delta_{\text{mix}}$ values are lower than that of the 2+1+1 flavor HISQ ensemble with the S(1) gauge action.
@@ IV Summary
Our central finding is the observation of a favorable $\mathcal{O}(a^{4})$ scaling for $\Delta_{\text{mix}}$ when valence overlap or clover fermions are used on gauge ensembles generated with 2+1+1 flavor HISQ sea fermions. This result is particularly significant as it was achieved using the same tadpole-improved Symanzik gauge action previously associated with a sizable $\Delta_{\text{mix}}$ in CLQCD clover fermion ensembles Zhao:2022ooq. This comparison directly verifies that the use of a dynamical fermion action with chiral symmetry is the dominant factor in suppressing mixed-action artifacts, irrespective of the gauge action. Furthermore, by comparing different action combinations at a fixed lattice spacing, we have quantified additional, secondary influences on $\Delta_{\text{mix}}$. Our results show that the larger correction from rectangle loops in the gauge action can provide an additional suppression of $\Delta_{\text{mix}}$, while the effect of dynamical charm quarks is statistically insignificant.
@@ IV.1 Autocorrelation Analysis
Figure 5: Autocorrelation analysis across multiple lattice ensembles, depicting the bin-size dependence of variance for four fundamental quantities: the plaquette operator, topological charge, pion mass, and $\eta_{c}$ meson mass. The notation $\{\mathrm{Traj.\ }n_{\mathrm{min}},-n_{\mathrm{max}},\mathrm{step\ }\Delta n\}$ accompanying each observable specifies the minimal and maximal trajectory indices, along with the stride parameter employed for configuration subsampling to ensure statistical independence in our analysis.